OpenStudy (anonymous):

Solve the system using matrices. -3x - y - 2z = -38 9x + 9z = 117 2y + z = 16

OpenStudy (anonymous):

Using matrices? How?$$\begin{bmatrix}-3&-1&-2\\9&0&9\\0&2&1\end{bmatrix}\begin{bmatrix}x\\y\\z\end{bmatrix}=\begin{bmatrix}-38\\117\\16\end{bmatrix}$$ Recognizing our coefficient matrix is invertible, you could just use its inverse to solve the system. You could also use elimination, of course -- either Gaussian or Gauss-Jordan. You could even use Cramer's rule.

OpenStudy (anonymous):

Thank you o: